OpenStudy (kylietipton1219999999999999999):

Point A is located at (0, 4), and point C is located at (−3, 5). Find the x value for point B that is located one over four the distance from point A to point C.

OpenStudy (kylietipton1219999999999999999):

A -0.25 B -0.5 C -0.75 D -1

OpenStudy (sshayer):

\[\frac{ 1 }{ 4 }::1:4\] sum of ratios=1+4=5